OpenStudy (anonymous):

What is the solution of the equation? 2 = 6p -8 -5p A. –10 B. –6 C. 2 D. 10

OpenStudy (nubeer):

2 = 6p - 8 -5p generally u have to make "P" alone in this case.. first add or subtract the terms with P.. can u?

OpenStudy (anonymous):

Honestly I dont know how to do these problems all I know is that its not a

OpenStudy (nubeer):

ok just look the terms with p.. its 6p and -5p agreed?

OpenStudy (anonymous):

Yes

OpenStudy (nubeer):

ok first subtract these terms it will leave. 2= 6p-5p-8 2 = p -8.. agreed?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then do I divide ,8 by 2

OpenStudy (anonymous):

-8

OpenStudy (nubeer):

no there is a sign so u dont have to divide.. add +8 on both sides..

OpenStudy (anonymous):

-6?

OpenStudy (nubeer):

hmm noo.. 2= p-8 right? now add +8 on both sides 2+8 = p -8 +8 10 = p.. get it?

OpenStudy (nubeer):

@jennyjewell25 do u get it? if not i can explain again.

OpenStudy (anonymous):

I got it thank you

OpenStudy (nubeer):

ok you're welcome..